Kyky232:

Pyramid A is a square pyramid with a base side length of 18 inches and a height of 9 inches. Pyramid B has a volume of 3,136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percentage. Provide an explanation and proof

Kyky232:

@snowflake0531

Florisalreadytaken:

hmm so there are a couple of steps we have to do. 1) find the area od the square at the borrom of the square pyramid \[ ( B=18 \times 18 ) \] then use the formula to find the volume of the whole pyramid: \[ V=\frac{B \times h}{3} \] \[ B= \text{look above}\] \[ h=9 \] do the math, and after you find the volume of the pyamid sivide the volume of pyramid A with pyramid B: \[ \frac{volume of pyramid A}{3.136} \]

Kyky232:

972 inches ^3 is the volume of the pyramid

Florisalreadytaken:

yes! now do the 2nd part

Kyky232:

309.948979592

Florisalreadytaken:

lets round that to 309.95 so pyramid A, is 309.95 times bigger than pyamid b how can we express that number as a percentage?

Kyky232:

i dont know :(

Florisalreadytaken:

we multiply that number by 100% so: \[ 309.95 \times 100\%=??\% \]

Florisalreadytaken:

\[309\cancel{.95} \times 1 \cancel{00}\%=30995\%\] so, pyramid A, is 30995% bigger than pyamid b.

surjithayer:

\[volume~of~pyramid~A=\frac{ base~area \times height }{ 3 }=\frac{ 18\times 18\times 9 }{3 }\] =972 inch ^3 volume of pyramid B=3,136 inch ^3 \[\frac{ volume ~of~pyramid~B }{ volume~of~pyramid~A}=\frac{ 3136 }{ 972 }\approx 3.2263374\] \[volume~of~pyramid~B ~is ~322.63 \% more~than~volume~of~pyramid~A\]

Florisalreadytaken:

wrong...

surjithayer:

it is not 3.136 but it is 3,136 means 3136

AZ:

I received a tag from this post but that reply was deleted. I do agree with surjithayer's work. You're supposed to divide Volume B / Volume A That's how you'll know how many times larger the second pyramid is from the first pyramid. And then you multiply by 100% It is also 3,136 and not 3.136 as surjithayer already pointed out.

Florisalreadytaken:

oh wow -- i somewhy had the idea of doing it the opposite way.